OpenStudy (anonymous):

you have a cable plan that cost 39 a month for a basic plan plus a movie channel your friend has the same basic plan plus two movie channels for 45.50 what is the basic plan charge they both pay

OpenStudy (anonymous):

Let x = cost of basic plan Let y = cost of movie channel We can create the following system of equations 39 = x + y 45.5 = x + 2y We can solve this system of equations using elimination by subtracting the first equation from the second equation. 45.5 = x + 2y -39 = -x - y -------------- 6.5 = 0 + y 6.5 = y So the cost of the movie channel is $6.50. We can back substitute (replace) y in the first original equation with 6.5 and solve for x. 39 = x + y 39 = x + (6.5) 39 - 6.5 = x + 6.5 - 6.5 32.5 = x + 0 32.5 = x The cost of the basic plan is $32.50 a month.